The perimeter of the rectangle below is 76 units. Find the length of side AD.
Write your answer without variables.
AD: 3x+2, AB: 4x+1

Answer by ikleyn(52785)

.

    (3x+2) + (4x+1) = 76/2 = 38   (half of the perimeter)


    7x + 3 = 38

    7x     = 38 - 3 = 35

     x              = 35/7 = 5 units.

    AD = 3x+2 = 3*5+2 = 17 units.      ANSWER

Solved.